# DoRA: Weight-Decomposed Low-Rank Adaptation

### v5

@@ 6.2 Text-to-Image Generation
Recently, as diffusion models have expanded in size, LoRA has become a popular method for efficiently fine-tuning large stable diffusion models. In this section, we aim to explore whether DoRA’s advantages over LoRA extend to the task of text-to-image generation. We follow the training pipeline of DreamBooth (Ruiz et al., 2023) for fine-tuning SDXL (Podell et al., 2023), utilizing the advanced training scripts developed by HuggingFace. The hyperparameter settings for LoRA and DoRA are kept the same, and we fine-tune the model using two challenging datasets: 3D icons and Lego sets. The sample seeds for generating the images are kept the same for LoRA and DoRA for fair comparison. The generated images are shown in Figure 8 and 9 in the appendix. The results indicate that DoRA achieves significantly better personalization than LoRA when using the same training settings, and more accurately reflects the training targets. For example, in Figure 8, the first sub-figure of DoRA’s output features a unique round square around the image, which is a feature common to all the training targets. In contrast, this feature is absent in all the LoRA outputs. A similar observation could be found with the Lego training targets, where only the DoRA outputs consistently incorporate the Lego logo in the generated images.
@@ A.3 Hyperparameters
A.4 Visual instruction tuning evaluation result
@@
A.5 Instruction-tuning evaluation examples
@@
A.6 Robustness towards different numbers of instruction tunning examples
@@
A.7 Robustness of DoRA towards different rank
@@
A.8 Text-to-Image Generation
Figures 8 and 9 show the images produced by SDXL fine-tuned with  DoRAand LoRA via DreamBooth (Ruiz et al., 2023) personalization techniques on two distinct training sets: 3D Icon [^1] https://huggingface.co/datasets/linoyts/3d_icon and Lego [^2] https://huggingface.co/datasets/merve/lego_sets_latest. The results reveal that DoRA can achieve considerably better personalization than LoRA with identical training configurations, more closely matching the training target.

### v6

@@ 5.1 Commonsense Reasoning
Figure 3: Magnitude (a) and direction (b) difference of LoRA/DoRA and the pre-trained weight of the query matrices across different layers.
Additionally, in previous sections, we hypothesize that a negative correlation between the magnitude update and directional update is more optimal than a positive correlation. This is because pre-trained weights already contain substantial knowledge suitable for downstream tasks, and a larger magnitude or direction alteration alone is sufficient for downstream adaptation. To further validate our hypothesis, we used LLaMA2-7B fine-tuned with DoRA/LoRA on commonsense reasoning datasets as a case study. We visualized the magnitude ($\Delta M$) and directional difference ($\Delta D$) between the DoRA/LoRA weights and the pre-trained model weights across different modules and layers. In Figure 3 (a) and (b), we observe that the DoRA fine-tuned weights show less deviation from the pre-trained weights in both magnitude and direction, while the differences for the LoRA fine-tuned weights are significantly larger. Coupled with the experimental results that DoRA significantly outperforms LoRA, we can conclude that our earlier hypothesis is valid: a robust foundation model does not require significant alterations for effective downstream adaptation and having the ability to perform more fine-grained magnitude and directional update explains the superiority of DoRA over LoRA. We leave the visualization of the value and key weight matrices in the appendix.
@@ 6.2 Text-to-Image Generation
Recently, as diffusion models have expanded in size, LoRA has become a popular method for efficiently fine-tuning large stable diffusion models. In this section, we aim to explore whether DoRA’s advantages over LoRA extend to the task of text-to-image generation. We follow the training pipeline of DreamBooth (Ruiz et al., 2023) for fine-tuning SDXL (Podell et al., 2023), utilizing the advanced training scripts developed by HuggingFace. The hyperparameter settings for LoRA and DoRA are kept the same, and we fine-tune the model using two challenging datasets: 3D icons and Lego sets. The sample seeds for generating the images are kept the same for LoRA and DoRA for fair comparison. The generated images are shown in Figure 10 and 11 in the appendix. The results indicate that DoRA achieves significantly better personalization than LoRA when using the same training settings, and more accurately reflects the training targets. For example, in Figure 10, the first sub-figure of DoRA’s output features a unique round square around the image, which is a feature common to all the training targets. In contrast, this feature is absent in all the LoRA outputs. A similar observation could be found with the Lego training targets, where only the DoRA outputs consistently incorporate the Lego logo in the generated images.
@@ A.3 Hyperparameters
A.4 Magnitude and Direction difference between DoRA/LoRA fine-tuned weight and the pre-triained weight of LLaMA2-7B for the commonsesne reasoning tasks
Figure 8 depicts the magnitude and direction differences in the weights of the query, key, and value matrices between LoRA/DoRA fine-tuned models and the pre-trained model across various layers of LLaMA2-7B for the commonsense reasoning tasks. The figure shows that the DoRA fine-tuned weights deviate less from the pre-trained weights in both magnitude and direction, supporting our hypothesis that a robust foundation model does not need substantial changes for effective downstream adaptation.
Figure 8: Magnitude and direction differences in the weights of the query, key, and value matrices between LoRA/DoRA fine-tuned models and the pre-trained model across various layers of LLaMA2-7B for the commonsense reasoning tasks.
A.5 Visual instruction tuning evaluation result
@@
A.6 Instruction-tuning evaluation examples
@@
A.7 Robustness towards different numbers of instruction tunning examples
@@
A.8 Robustness of DoRA towards different rank
@@
A.9 Text-to-Image Generation
Figures 10 and 11 show the images produced by SDXL fine-tuned with  DoRAand LoRA via DreamBooth (Ruiz et al., 2023) personalization techniques on two distinct training sets: 3D Icon [^1] https://huggingface.co/datasets/linoyts/3d_icon and Lego [^2] https://huggingface.co/datasets/merve/lego_sets_latest. The results reveal that DoRA can achieve considerably better personalization than LoRA with identical training configurations, more closely matching the training target.